Complete the balance sheet and sales information using the following financial data:

Total assets turnover: 1.2x
Days sales outstanding: 42 daysa
Inventory turnover ratio: 7x
Fixed assets turnover: 3.5x
Current ratio: 2.3x
Gross profit margin on sales: (Sales - Cost of goods sold)/Sales = 35%
aCalculation is based on a 365-day year. Do not round intermediate calculations. Round your answer to the nearest cent.

Balance Sheet
Cash $   Current liabilities $  
Accounts receivable    Long-term debt 87,500
Inventories    Common stock   
Fixed assets    Retained earnings 87,500
Total assets $350,000 Total liabilities and equity $  
Sales $   Cost of goods sold $  
Answer #1

Total asset turnover = 1.2

Sales/total assets = 1.2

Sales = 1.2*350,000 =$420,000

Cost of goods sold =420,000*65% =$273,000

Days sales outstanding = 42

365*accounts receivables/sales =42

Accounts receivables = 42*420,000/365

= 48,328.77

Inventory turnover ratio = 7

Cost of goods sold/inventory = 7

Inventory = 273000/7 =$39,000

Fixed assets turnover = 3.5

Sales/fixed assets = 3.5

Fixed assets = 420,000/3.5

=$120,000

Cash = 350,000-120,000-48,328.77-39,000

=$142,671.2

Current ratio = 2.3

Current assets/current liabilities = 2.3

Current liabilities = 230,000/2.3

=$100,000

Common stock =$75,000

Total liabilities and equity = 350,000